### v1

# Anisotropic flow fluctuation as a possible signature of clustered nuclear geometry in O–O collisions at the Large Hadron Collider

Abstract
Nuclei having $4n$ number of nucleons are theorized to possess clusters of $\alpha$ particles (4He nucleus). The Oxygen nucleus (16O) is a doubly magic nucleus, where the presence of an $\alpha$-clustered nuclear structure grants additional nuclear stability. In this study, we exploit the anisotropic flow coefficients to discern the effects of an $\alpha$-clustered nuclear geometry w.r.t. a Woods-Saxon nuclear distribution in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using a hybrid of IP-Glasma + MUSIC + iSS + UrQMD models. In addition, we use the multi-particle cumulants method to measure anisotropic flow coefficients, such as elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), as a function of collision centrality. Anisotropic flow fluctuations, which are expected to be larger in small collision systems, are also studied for the first time in O–O collisions. It is found that an $\alpha$-clustered nuclear distribution gives rise to an enhanced value of $v_{2}$ and $v_{3}$ towards the highest multiplicity classes. Consequently, a rise in $v_{3}/v_{2}$ is also observed for the (0-10)centrality class. Further, for $\alpha$-clustered O–O collisions, fluctuations of $v_{2}$ are larger for the most central collisions, which decrease towards the mid-central collisions. In contrast, for a Woods-Saxon 16O nucleus, $v_{2}$ fluctuations show an opposite behavior with centrality. This study, when confronted to experimental data may reveal the importance of nuclear density profile on the discussed observables.
I Introduction
The primary goal of relativistic heavy-ion collisions at the RHIC and the LHC is to create a deconfined thermalised medium of partons, also known as Quark Gluon Plasma (QGP), which is, by our understanding, have existed a few microseconds after the Big Bang. However, due to the very low lifespan of QGP, it can not be observed directly in experiments. Instead, several indirect signatures from experimentally measurable observables can signify the presence of such a deconfined strongly interacting matter in heavy-ion collisions. One such observable is azimuthal anisotropy, which characterises the collective behaviour of the medium formed in heavy-ion collisions. It is quantified in terms of the Fourier expansion coefficients of azimuthal distribution of final state charged hadrons, as follows [1]:
$\frac{dN}{d\phi}\propto\Bigg{(}1+2\sum_{n=1}^{\infty}v_{n}\cos\left[n(\phi-\psi_{n})\right]\Bigg{)}\ .$ (1)
Here, $\phi$ is the azimuthal angle, $\psi_{n}$ stands for the $n$th order symmetry plane angle according to the give $v_{n}$ coefficients of the Fourier expansion. Here, $v_{2}$ quantifies the elliptic flow, $v_{3}$ gives the estimates for triangular flow, and so on. Both $v_{2}$ and $v_{3}$ depend upon the equation of state and transport properties of the QCD medium formed during heavy-ion collisions. In addition, studies based on hydrodynamical calculations show that the flow coefficients are affected strongly by a change in the ratio of shear viscosity to entropy density ($\eta/s$[2, 3]. Moreover, the sensitivity of flow coefficients to $\eta/s$ is large for higher-order harmonics, i.e., $v_{3}$ is more sensitive to $\eta/s$ as compared to $v_{2}$ [3]. In several MC studies, $v_{2}$ and $v_{3}$ are found to be associated with the initial eccentricities of the participating nucleons, where the values of anisotropic flow coefficients can change by modifying the nuclear distribution [6, 4, 5]. In addition, in Xe–Xe collisions, an observation of a large value of $v_{2}$ in most central collisions, as compared to that observed in Pb–Pb collisions, is anticipated due to a deformed nuclear structure of Xe nucleus [7, 8, 10, 9]. In a similar line, the nuclear deformation of Uranium, being a prolate shape nucleus is reflected in a higher $v_{2}$, when compared to that is observed in the most central collisions of Au nuclei, which is close to a spherical nucleus [11]. Similarly, one notices that the values of $v_{2}$ and $v_{3}$ vary with the change in the deformation parameters of Uranium in U–U collisions, and the effect is enhanced for $v_{2}$, as compared to $v_{3}$ [4, 5, 12].
Figure 1: Pictorial representation of an $\alpha$-clustered (left) and non-clustered (right) distribution of nucleons inside the oxygen nucleus.
Elliptic flow is one of the observables that can probe the initial state of heavy-ion collisions [13]. However, the value of elliptic flow fluctuates from one event to another, which reflects the fluctuation in the initial energy density and event-by-event participant nucleon distribution before QGP is formed [14]. These fluctuations in the initial density profile of the participating nucleons can lead to fluctuations in the even harmonic flow coefficients and generate odd flow harmonics [14]. Further, the fluctuations in the initial energy density can develop symmetry planes of different orders in various kinematic regions [15, 16, 17]. In other words, the initial density fluctuations can develop transverse momentum ($p_{\rm T}$) and pseudorapidity ($\eta$) dependence of $\psi_{n}$. Therefore, to comprehend the event-by-event initial density fluctuations and their impact on medium evolution, the study of anisotropic flow fluctuations becomes crucial. In both experimental and theoretical frontiers, a few studies on anisotropic flow fluctuations have been performed in heavy-ion collisions [18, 19, 20, 21]. The Pb–Pb collisions at the LHC energies hint at non-Gaussian initial density fluctuations, and consequently, it led the heavy-ion physicists to constrain their probability distribution functions [22, 23]. However, the studies of anisotropic flow fluctuations in small collision systems are not yet well understood. Moreover, one may conceive that for a small number of participants, the contribution to fluctuations can increase. Finally, if the evolution of anisotropic flow fluctuations has more than one contributor, then the results can vary significantly from the expectations.
In addition to the collective behaviour of the QGP, several other signatures for the presence of a QGP medium have been observed in heavy-ion collisions. For the measurement of such signatures, the baseline measurements are obtained from proton-proton ($pp$) collisions, where the formation of a QCD medium is not anticipated. However, recent observations of strangeness enhancement [24], ridge-like structure [26, 25], radial flow effects [27, 28, 29], etc., in high multiplicity $pp$ collisions have highlighted the significance of studying the small collision systems from different aspects. Recently, LHC and RHIC have a plan to inject Oxygen nuclei (16O) to perform p-O and O–O collisions [30, 31], which can bridge the multiplicity gap between $pp$, p-Pb and Pb–Pb collisions and may provide insight into the possible formation of QGP droplets in small systems. Furthermore, Oxygen is a stable nucleus having a double magic number and is theorized to possess clusters of $\alpha$ (4He) particles, where the $\alpha$-particles arrange themselves in the corners of a tetrahedron [32, 33, 34, 35, 36, 37, 38]. Figure 1 shows a pictorial representation of the nucleons inside an oxygen (16O) nucleus for an $\alpha$-clustered structure (left) and non-clustered distribution of nucleons (right). In recent years, there have been several studies performed to understand the O–O collisions both at RHIC and LHC energies [44, 45, 43, 39, 40, 41, 6, 42, 46]. A few of them investigate the use of hydrodynamical models [47, 48, 49], and Glauber model calculations [39, 40, 41]. Some intend to study the parton energy loss and jet quenching effects in O–O collisions [50, 51]. Interestingly, there have been a few phenomenological studies that aim to establish the effect of the presence of an $\alpha$-cluster nucleus in the final state particle production [44, 45, 43, 53, 54, 55, 56, 6, 52, 57, 58, 59]. A few of these studies employ O–O collisions, while others employ O-Au, C-Au, O-Pb, and Ne-Pb collisions. Nevertheless, the presence of a clustered nuclear structure of 16O is reflected in final state azimuthal anisotropy when compared to a Woods-Saxon nuclear density profile. In Ref. [52], using a multi-phase transport model, authors have shown that an away side broadening is observed in the two-particle correlation function for an $\alpha$-clustered particle. In addition, an increased value of $v_{3}$, and $v_{3}/v_{2}$ for most central O–O collisions are also reported for an $\alpha$-clustered structure. A similar observation was reported in Ref. [44], where authors study the anisotropic flow in C+Au collisions with a clustered structure of 12C nucleus. However, so far, no studies have shown the effect of change in nuclear density profile on event-by-event flow fluctuations. In this paper, for the first time, we report the behaviour of azimuthal anisotropy and flow fluctuations in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using a hydrodynamically expanding system. For this study, we use a hybrid of IPGlasma initial condition, MUSIC hydrodynamics, iSS as a particlisation and UrQMD as an afterburner for the hadronic cascade. Here, we compared results from an $\alpha$-clustered nuclear distribution with a Woods-Saxon nuclear distribution.
This paper is organized as follows. We start with a brief introduction and motivation to this study in Section I. We briefly discuss the event generation and methodology in Section II. Then, we present our findings and discuss the results in Section III. Finally, we summarise our findings with a brief outlook in Section IV.
II Event Generation and Methodology
In this section, we briefly introduce the hybrid simulation framework used to generate events, including a discussion on the multi-particle cumulant method for the anisotropic flow estimation.
II.1 Simulation framework
In this work, the simulation framework implements the impact-parameter-dependent Glasma (IP-Glasma) initial conditions, MUSIC relativistic viscous hydrodynamics, particle sampling using the iSS package, and performs hadronic cascade through the UrQMD transport model. This model provides a good description of the particle production and flow in heavy-ion collisions [60]. The key aspects of these models are also discussed briefly below, including the parameters and settings necessary to reproduce the results reported in this study.
II.1.1 IP-Glasma: Initial condition
The IP-Glasma model is used to describe the dynamics of the gluon fields during the collisions of nuclear matter at relativistic speeds [61, 62]. It provides the initial energy-momentum tensor of the classical Yang-Mills (CYM) color Glasma fields ($T_{\rm YM}^{\mu\nu}$) which can be further evolved using relativistic hydrodynamics. This model is based on the color glass condensate effective theory [63, 64] and uses a classical description of gluon production. This model includes fluctuations in the distributions of nucleons in the nuclear wave functions and also fluctuations in the color charge distributions inside the nucleons. The color charge fluctuations are modeled via the impact-parameter-dependent dipole saturation model (IP-Sat) [65, 66]. For the sub-nucleonic fluctuations, three hot spots are introduced in the proton thickness function, and each hot spot is parametrized using a Gaussian distribution in the transverse plane. The spatial positions of the nucleons inside the colliding nuclei are obtained from the Woods-Saxon distribution function as given below,
$\rho(r)=\rho_{0}\frac{1+w\left(\frac{r}{r_{0}}\right)^{2}}{1+\exp\left(\frac{r-r_{0}}{a}\right)}\ .$ (2)
Here, $\rho(r)$ is the nuclear charge density at a radial distance $r$ from the center of the nucleus. $r_{0}$, $w$, and $a$ are the mean radius of the nucleus, the nuclear deformation parameter, and nuclear skin depth, respectively. For 16O nucleus, $r_{0}=2.608$ fm, $a=0.513$ fm, $w=-0.051$ [67]. In addition, the exotic $\alpha$–cluster geometry for the 16O nucleus has also been implemented in this study. Four such $\alpha$–particles ($2p,~{}2n$) are placed at the vertices of a regular tetrahedron with side length 3.42 fm, which makes the root mean square (RMS) radius of 16O nucleus to 2.699 fm [6, 42, 43]. The spatial positions of the nucleons inside an $\alpha$–particle are sampled using the Woods-Saxon distribution given in Eq. (2), with the parameters chosen for the 4He nucleus, which are $r_{0}=0.964$ fm, $w=0.517$ and $a=0.322$ fm. This gives the RMS radius of 1.676 fm for the $\alpha$–particle. A minimum separation distance of 0.4 fm between the nucleons is imposed. Finally, to randomize each nucleus, the entire tetrahedral configuration is rotated through its center randomly in the polar and azimuthal coordinates before its initialization in IP-Glasma. Figure 2 shows the nucleon distribution of 16O nuclei in the x-y plane following $\alpha$-cluster (left) and Woods-Saxon (right) nuclear density profiles. In Fig. 2, for the $\alpha$-cluster nuclear density profile of 16O nuclei, the tetrahedral configuration is not rotated for better visualization of the clustered geometry.
Figure 2: The nucleon distribution of 16O nuclei in the x-y plane following $\alpha$-cluster (left) and Woods-Saxon (right) nuclear density profiles plotted with one million nuclei in each case.
Once the nucleon centers are obtained, the color charges are sampled using the IP-Sat/McLerran-Venugopalan model [64, 68]. Now, color currents arise due to the motion of color charges present inside the colliding nuclei along the beam direction. These currents act as the sources in the Yang-Mills equation, $[D_{\mu},F^{\mu\nu}]=J^{\nu}$. The initial condition at $\tau=0$ is given by the solution of CYM equations in the Schwinger gauge. For the subsequent evolution of the equations of motion for Glasma fields, a lattice formulation has been used, which was first introduced in Ref. [69]. These equations of motion are solved using a leapfrog algorithm on a 2D lattice. For our case, we use the lattice size, $L=14$, and lattice spacing $a=0.02$ fm [49]. Finally, all the component of the energy-momentum tensor ($T_{\rm YM}^{\mu\nu}$) for the Yang-Mills system is computed at $\tau_{\rm switch}=0.4$ fm, and the initial-state configuration is transferred to the hydrodynamic simulation.
II.1.2 MUSIC: Hydrodynamics
Under the assumption of system under local thermal equilibrium, the initial energy-momentum tensor obtained from the IP-Glasma model at thermalization time $\tau_{0}=0.4$ fm is evolved using the MUSIC framework of relativistic hydrodynamics following the conservation law $\partial_{\mu}T^{\mu\nu}=0$ [70, 71, 72]. The energy-momentum tensor is composed of both ideal and dissipative parts, which are combined to give, $T^{\mu\nu}=eu^{\mu}u^{\nu}-(P+\Pi)\Delta^{\mu\nu}+\pi^{\mu\nu}$. Here, $e$ is the energy density, $u^{\mu}$ is the fluid flow four-velocity, $P$ is the thermodynamic pressure, $\Pi$ is the bulk viscous pressure, and $\pi^{\mu\nu}$ is the shear viscous tensor. Again, the projection operator is defined as $\Delta^{\mu\nu}=g^{\mu\nu}-u^{\mu}u^{\nu}$. In this study, we use the boost invariant picture of hydrodynamics and use the equation-of-state parametrization "s95p-v1.2", which is obtained from the interface of lattice data and hadron resonance gas [73]. We use a fixed $\eta/s=0.12$, and a temperature dependent $\zeta/s(T)$ as described in Ref. [49]. In MUSIC, the hydrodynamic equations are solved using the Kurganov-Tadmor (KT) algorithm [74, 75].
II.1.3 iSS: Particle sampler
The fluid-dynamic description is finally switched to the particle description when the local energy density drops to the switching energy density, $e_{\rm sw}=0.18$ GeV/fm3 [49]. At this point, the freeze-out hypersurface from MUSIC is fed to the iSS particle sampler [76, 77], which follows the usual Cooper-Frye formalism for computing the particle spectra using equilibrium distribution with both shear and bulk viscous correction terms [78, 79, 49]. Each MUSIC freeze-out hypersurface can be over-sampled using the iSS particle sampler to increase event statistics and save computation time. So, we have sampled 200 events from each MUSIC output hypersurface.
II.1.4 UrQMD: Hadronic cascade
Finally, the produced particles are evolved using the UrQMD microscopic transport model [80, 81]. This model handles both elastic and inelastic processes of hadronic scatterings and resonance decay with particles of masses up to 2.25 GeV. UrQMD simulates a more realistic description of the final-stage hadron evolution and performs a dynamical-freezeout for different particle species. All parameters in the UrQMD model are set to their default values. The output of UrQMD includes final-state particle four momenta and particle identification (PID) code which can be written to disk for further analysis.
II.2 Multiparticle cumulant method
Anisotropic flow can be expressed as the coefficients of Fourier expansion, as shown in Eq. (1). One can estimate the anisotropic flow coefficients using the following equation [82, 83, 84]:
$v_{n}=\langle\cos[n(\phi-\psi_{n})]\rangle\ .$ (3)
[…]
$Q_{n}=\sum_{j=1}^{M}e^{in\phi_{j}},$ (4)
where $M$ is the multiplicity of the event, and $\phi$ is the azimuthal angle of the charged hadrons. One can obtain a single event average two- and four-particle correlation function using the following expressions.
$\langle 2\rangle$ $=\frac{|Q_{n}|^{2}-M}{M(M-1)},$ (5)
$\langle 4\rangle$ $=\frac{|Q_{n}|^{4}+|Q_{2n}|^{2}-2\cdot{\rm{Re}}[Q_{2n}Q_{n}^{*}Q_{n}^{*}]}{M(M-1)(M-2)(M-3)}$
$~{}~{}-2\frac{2(M-2)\cdot|Q_{n}|^{2}-M(M-3)}{M(M-1)(M-2)},$
One can obtain the event-average correlation function using the following expression.
$\langle\langle 2\rangle\rangle$ $=\frac{\sum_{i=1}^{N_{\rm ev}}(W_{\langle 2\rangle})_{i}\langle 2\rangle_{i}}{\sum_{i=1}^{N_{\rm ev}}(W_{\langle 2\rangle})_{i}},$ (6)
$\langle\langle 4\rangle\rangle$ $=\frac{\sum_{i=1}^{N_{\rm ev}}(W_{\langle 4\rangle})_{i}\langle 4\rangle_{i}}{\sum_{i=1}^{N_{\rm ev}}(W_{\langle 4\rangle})_{i}}.$
[…]
$W_{\langle 2\rangle}$ $=M(M-1),$ (7)
$W_{\langle 4\rangle}$ $=M(M-1)(M-2)(M-3).$
Consequently, the two- and four-particle cumulants can be obtained as:
$c_{n}\{2\}$ $=\langle\langle 2\rangle\rangle,$ (8)
$c_{n}\{4\}$ $=\langle\langle 4\rangle\rangle-2\cdot\langle\langle 2\rangle\rangle^{2},$
Now, the reference flow of the particles can be estimated using the following expression.
$v_{n}\{2\}$ $=\sqrt{c_{n}\{2\}},$ (9)
$v_{n}\{4\}$ =$4]{-c_{n}\{4\}}.$
To estimate the differential flow of the particles of interest (POIs), the $p_{n}$ and $q_{n}$ vectors for specific kinematic range and/or for specific hadron species are defined as follows:
$p_{n}$ $=\sum_{j=1}^{m_{p}}e^{in\phi_{j}},$ (10)
$q_{n}$ $=\sum_{j=1}^{m_{q}}e^{in\phi_{j}},$
where $m_{p}$ is the total number of particles labeled as POIs, and $m_{q}$ is the total number of particles tagged as both reference flow particles (RFPs) and POIs. RFPs serve as a reference frame for POIs to quantify the collective motion of the system and help establish the orientation of the event plane, which is crucial for determining the anisotropic flow coefficients. The single-event average differential two- and four-particle azimuthal correlation functions are calculated as:
$\langle 2^{{}^{\prime}}\rangle$ $=\frac{p_{n}Q_{n}^{*}-m_{q}}{m_{p}M-m_{q}},$ (11)
$\langle 4^{{}^{\prime}}\rangle$ $=\left[p_{n}Q_{n}Q_{n}^{*}Q_{n}^{*}-q_{2n}Q_{n}^{*}Q_{n}^{*}-p_{n}Q_{n}Q_{2n}^{*}\right.$
$-2\cdot Mp_{n}Q_{n}^{*}-2\cdot m_{q}|Q_{n}|^{2}+7\cdot q_{n}Q_{n}^{*}-Q_{n}q_{n}^{*}$
$\left.+q_{2n}Q_{2n}^{*}+2\cdot p_{n}Q_{n}^{*}+2\cdot m_{q}M-6\cdot m_{q}\right]$
$/\left[(m_{p}M-3m_{q})(M-1)(M-2)(M-3)\right]$
Similarly, one obtains the event-average differential azimuthal correlations using the following expressions.
$\langle\langle 2^{{}^{\prime}}\rangle\rangle$ $=\frac{\sum_{i=1}^{N_{\rm ev}}(w_{\langle 2^{{}^{\prime}}\rangle})_{i}\langle 2^{{}^{\prime}}\rangle_{i}}{\sum_{i=1}^{N_{\rm ev}}(w_{\langle 2^{{}^{\prime}}\rangle})_{i}},$ (12)
$\langle\langle 4^{{}^{\prime}}\rangle\rangle$ $=\frac{\sum_{i=1}^{N_{\rm ev}}(w_{\langle 4^{{}^{\prime}}\rangle})_{i}\langle 4^{{}^{\prime}}\rangle_{i}}{\sum_{i=1}^{N_{\rm ev}}(w_{\langle 4^{{}^{\prime}}\rangle})_{i}}.$
Here, $w_{\langle 2^{{}^{\prime}}\rangle}$ and $w_{\langle 4^{{}^{\prime}}\rangle}$ are the weights corresponding to two- and four-particle cumulants, given by:
$w_{\langle 2^{{}^{\prime}}\rangle}$ $=m_{p}M-m_{q},$ (13)
$w_{\langle 4^{{}^{\prime}}\rangle}$ $=(m_{p}M-3m_{q})(M-1)(M-2)(M-3).$
Thus, one finds the two- and four-particle differential cumulants using the following equations:
$d_{n}\{2\}$ $=\langle\langle 2^{{}^{\prime}}\rangle\rangle,$ (14)
$d_{n}\{4\}$ $=\langle\langle 4^{{}^{\prime}}\rangle\rangle-2\langle\langle 2^{{}^{\prime}}\rangle\rangle\langle\langle 2\rangle\rangle.$
Finally, one can calculate the differential flow $v_{2}(p_{\rm T})$ using two- and four-particle correlations as follows:
$v_{n}\{2\}(p_{\rm T})$ $=\frac{d_{n}\{2\}}{\sqrt{c_{n}\{2\}}},$ (15)
$v_{n}\{4\}(p_{\rm T})$ $=-\frac{d_{n}\{4\}}{\sqrt[4]{(-c_{n}\{4\})^{3}}}.$
Figure 3: Pictorial representation of the pseudorapidity regions $A$ and $B$ separated with a pseudorapidity gap ($\Delta\eta$).
Unfortunately, the $v_{n}$ obtained from the two-particle Q-Cumulant method may contain contributions from non-flow effects, which can be suppressed by appropriate kinematic cuts. One can introduce a pseudorapidity gap between the particles in the two-particle Q-cumulant method [91, 94, 93] to suppress the non-flow contributions. Accordingly, the whole event is divided into two sub-events [95, 96, 97], $A$ and $B$, which are separated by a $|\Delta\eta|$ gap, as depicted in Fig. 3. This modifies Eq. (5) to:
$\langle 2\rangle_{\Delta\eta}=\frac{Q_{n}^{A}\cdot Q_{n}^{B*}}{M_{A}\cdot M_{B}},$ (16)
where $Q_{n}^{A}$ and $Q_{n}^{B}$ are the flow vectors from the sub-event $A$ and $B$, respectively. $M_{A}$ and $M_{B}$ are the corresponding multiplicities. Thus, the two-particle Q-cumulant with a $|\Delta\eta|$ gap is given by:
$c_{n}\{2,|\Delta\eta|\}=\langle\langle 2\rangle\rangle_{\Delta\eta}.$ (17)
In calculations of differential flow with a pseudorapidity gap, there is no overlap of POIs and RFPs if we select RPs from one subevent and POIs from the other. Thus Eq. (11) can be modified to:
$\langle 2^{{}^{\prime}}\rangle_{\Delta\eta}=\frac{p_{n,A}Q_{n,B}^{*}}{m_{p,A}M_{B}},$ (18)
and we obtain the differential two-particle cumulant as:
$d_{n}\{2,|\Delta\eta|\}=\langle\langle 2^{{}^{\prime}}\rangle\rangle_{\Delta\eta}.$ (19)
Finally, the differential flow from the two-particle cumulant can be obtained by inserting the two-particle reference flow (with $\eta$ gap) into the differential two-particle cumulant:
$v_{n}\{2,|\Delta\eta|\}(p_{\rm T})=\frac{d_{n}\{2,|\Delta\eta|\}}{\sqrt{c_{n}\{2,|\Delta\eta|\}}}.$ (20)
In this paper, the elliptic and triangular flow coefficients are calculated using the above equations by setting $n=$ 2 and 3, respectively. Using Eqs. (9) and (20), one can obtain the mean and fluctuations of anisotropic flow using the following expressions.
$\langle v_{n}\rangle=\sqrt{\frac{v_{n}^{2}\{2\}+v_{n}^{2}\{4\}}{2}}$ (21)
$\sigma_{v_{n}}=\sqrt{\frac{v_{n}^{2}\{2\}-v_{n}^{2}\{4\}}{2}}$ (22)
Using Eqs. (21), and (22), one can obtain relative anisotropic flow fluctuation as follows.
$F(v_{n})=\frac{\sigma_{v_{n}}}{\langle v_{n}\rangle}$ (23)
In this study, we use all the charged hadrons within the pseudorapidity region, $|\eta|<2.5$, for the estimation of anisotropic flow coefficients using the two- and four-particle Q-cumulants method. We tag the charged hadrons within $|\eta|<2.5$ and $4.00.2<p_{\rm T}<$ 4.0 GeV/c as RFPs. We apply a pseudorapidity gap, $|\Delta\eta|>1.0$ in the two-subevent method, as shown in Fig. 3, to subtract the non-flow effects from two-particle Q-cumulants method. The statistical uncertainties are calculated using the relations shown in Ref. [98].
III Results and discussions
In this section, we start with a brief discussion of the multiplicity selection, followed by discussions on the results for flow coefficients. Finally, we discuss the centrality dependence of flow fluctuations.
Figure 4: Distribution of charged particle produced in the pseudorapidity acceptance of $-3.7<\eta<-1.7$ and $5.12.8<\eta<$ 5.1 (V0 Multiplicity or V0M) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon (upper) and $\alpha$-cluster (middle) nuclear density profiles. The lower panel shows the ratio of V0M distributions from Woods-Saxon to that of in $\alpha$-cluster nuclear profile.
Figure 4 shows the distribution of charged particles produced in the pseudorapidity region $-3.7<\eta<-1.7$ and $5.12.8<\eta<$ 5.1 (V0M) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for a Woods-Saxon (upper) and $\alpha$-cluster (lower) nuclear density profiles. The lower panel shows the ratio of normalized yields of V0M for Woods-Saxon and $\alpha$-cluster nuclear distributions. Also, different centrality classes are defined based on the percentile slices on the V0M distribution for the Woods-Saxon nuclear density profile. Here, for a fair comparison, we use the same cuts of V0M distribution for all the centrality classes for both the nuclear density profiles. The lowest central class, e.g., (0-10)%, refers to the most central collisions, while the highest centrality class implies a peripheral collision. One can refer to Fig. 8 in the Appendix .1 for the average value of the impact parameter corresponding to each centrality class for both Woods-Saxon and $\alpha$-cluster nuclear density profiles. In Fig. 4, one finds a higher yield for V0M for the O–O collisions having an $\alpha$-clustered nucleus. In addition, for a similar value of V0M, one finds a higher yield for the $\alpha$-clustered distribution on nucleons inside the oxygen nuclei for V0M $>300$, which can be seen in the bottom ratio plot. This signifies the compact nature of an $\alpha$-clustered oxygen nucleus, which is reflected in the most central O–O collisions.
Figure 5: Upper panel shows the elliptic flow and triangular flow using two-particle Q-cumulant method as a function of centrality class (%) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. The lower panel shows the ratio of $v_{n}\{2\}$ for Woods-Saxon to a $\alpha$-cluster nuclear density profile.
Figure 6: Upper panel shows the ratio $v_{3}\{2\}/v_{2}\{2\}$ as a function of centrality class (%) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. Lower panel shows the ratio of $v_{3}\{2\}/v_{2}\{2\}$ from Woods-Saxon to a $\alpha$-cluster nuclear density profile.
Figure 5 shows $v_{2}\{2\}$ and $v_{3}\{2\}$ with a pseudorapidity gap, $|\Delta\eta|>1.0$, as a function of centrality for Woods-Saxon and $\alpha$-cluster nuclear density profiles of 16O in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model. One notices a fair centrality dependence on both elliptic and triangular flow in both Woods-Saxon and $\alpha$-cluster nuclear density profiles. The values of $v_{2}\{2\}$ for the $\alpha$-cluster nuclear density profile are higher than the Woods-Saxon nuclear density profiles throughout the centrality classes except for the peripheral collisions, where the trend is reversed. Here, the value of elliptic flow is highest for the central collisions and decreases towards the peripheral collisions. In most central collisions, the observed high values of elliptic flow can be attributed to the initial eccentricity fluctuations. In contrast, towards the mid-central collisions, although the geometry of the overlap region plays a role, the presence of a lower number of participants makes the initial eccentricity completely transform into the final state elliptic flow of hadrons, thus lowering the value of elliptic flow. However, The observed behavior of elliptic flow as a function of collision centrality and nuclear density profiles assuming a hydrodynamical expansion vary significantly from that observed in a kinetic theory-based model, such as AMPT [6]. In contrast, we observe an interesting behavior for $v_{3}\{2\}$ as a function of collision centrality for both the nuclear density profiles. One finds that $v_{3}\{2\}$ decreases from most central to peripheral collisions, having similar slopes for the Woods-Saxon and $\alpha$-clustered nuclear density profiles within (20-90)centrality classes. This consistency in the triangular flow between two nuclear density profiles for the mentioned centrality class is consistent with the AMPT predictions, as shown in Ref. [6]. Interestingly, it can be observed from the lower panel of Figure 5, a sudden rise of $v_{3}\{2\}$ in the more central collisions for the $\alpha$-clustered in contrast to a Woods-Saxon density profile. These results are similar to the AMPT model calculations in Ref. [6]. We believe that the sudden rise in $v_{3}\{2\}$ for most central collisions of oxygen nuclei having an $\alpha$-clustered nuclear density profile can be attributed to the presence of an initial tetrahedral arrangement of four $\alpha$-particles inside the oxygen nucleus.
Figure 7: Collision centrality dependence of $v_{2}$, $\langle v_{2}\rangle$, $\sigma_{v_{2}}$, and $F(v_{2})$ in O–O collisions, from (a) to (d), respectively, at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles.
The upper panel of Fig. 6 shows the ratio $v_{3}\{2\}/v_{2}\{2\}$ as a function of collision centrality in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. The bottom panel shows the ratio of $v_{3}\{2\}/v_{2}\{2\}$ for Woods-Saxon to a $\alpha$-cluster nuclear density profile. Interestingly, when the oxygen nuclei having the Woods-Saxon nuclear distributions collide, their elliptic and triangular flow values follow a smooth dependence on collision centrality, as one can see in Fig. 5. Thus, the ratio $v_{3}\{2\}/v_{2}\{2\}$ also exhibits a smooth decrease from most central-to-peripheral collisions. However, one finds a notably sharp increase in $v_{3}\{2\}/v_{2}\{2\}$ for (0-5)centrality case for the $\alpha$-cluster case, followed by a smooth decrease in $v_{3}\{2\}/v_{2}\{2\}$ as a function of collision centrality. This reflects the presence of a tetrahedral arrangement of $\alpha$-clusters inside the oxygen nuclei giving rise to a large intrinsic triangular flow in the most central collisions. Although one finds $v_{2}\{2\}$ for $\alpha$-cluster density profile is larger compared to that of Woods-Saxon nuclear density profile, this tetrahedral structure raises $v_{3}\{2\}$ even higher for the most central collisions, which is also consistent with transport models such as AMPT, shown in Ref. [6]. Thus the observation of a sudden rise in tetrahedral $\alpha$-clustered case can lead to an observation of a spike in $v_{3}\{2\}/v_{2}\{2\}$ for most central collisions and consequently, the observation of which in O–O collisions may signify the presence of a $\alpha$-clustered nuclear geometry of 16O nuclei.
Figure 7 (a) shows the evolution of elliptic flow using two- and four-particle Q-cumulant method as a function of centrality for Woods-Saxon and $\alpha$-cluster nuclear density profiles in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model. One finds a small centrality dependence for $v_{2}\{2,|\Delta\eta|>1.0\}$ for both the nuclear density profiles within (0-30)centrality. In contrast, using the four-particle cumulants method, we observe a significant centrality dependence of elliptic flow. The value of $v_{2}\{4\}$ for the $\alpha$-cluster case increases from (0-10)to (10-20)%, and again decreases for (20-30)centrality. However, for the Woods-Saxon case, $v_{2}\{4\}$ maintains a consistent decreasing feature with an increase in the centrality. Another important point to note here is that $v_{2}\{2,|\Delta\eta|>1.0\}>v_{2}\{4\}$ for both the nuclear density profiles, which can consequently lead to real values of flow fluctuations. Figure 7 (b) shows the mean value of elliptic flow ($\langle v_{2}\rangle$), evaluated using Eq. (21), for Woods-Saxon and $\alpha$-cluster nuclear density profiles in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model. $\langle v_{2}\rangle$ is found to be larger for O–O collisions having a $\alpha$-cluster nuclear density profile as compared to a Woods-Saxon distribution. The centrality dependence of $\langle v_{2}\rangle$ for both the nuclear density profiles is similar to $v_{2}\{4\}$, as $v_{2}\{2,|\Delta\eta|>1.0\}$ is nearly independent of collision centrality in (0-30)centrality region. Figure 7 (c) shows a centrality dependence of elliptic flow fluctuation ($\sigma_{v_{2}}$) for Woods-Saxon and $\alpha$-cluster nuclear density profiles in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model. One finds that $\sigma_{v_{2}}$ for Woods-Saxon and $\alpha$-cluster nuclear density profiles have complementing centrality dependence, i.e., when going from most central to mid-central collisions, $\sigma_{v_{2}}$ for Woods-Saxon increases while for an $\alpha$-cluster nuclei it shows a decreasing trend. This makes one of the interesting distinctions between the two nuclear density profiles with the study of flow fluctuations. Although, in Pb–Pb collisions, due to the absence of elliptic geometry in the initial state, the elliptic flow in the most-central collisions has a large contribution from elliptic flow fluctuations; however, towards the mid-central collisions, the contribution of elliptic flow-fluctuations to elliptic flow decreases, as the elliptic geometry of the nuclear participants leads the elliptic flow contribution [93, 96, 97]. Interestingly, the $\alpha$-cluster case of O–O collisions shows a similarity in the evolution of elliptic flow fluctuations with the Pb–Pb collisions, which is clearly different in O–O collisions with a Woods-Saxon nuclear density profile. The source of this opposite behavior of $\sigma_{v_{2}}$ for both the nuclear density profiles can thus intrinsically be attributed to the difference in the nuclear geometry of the 16O nucleus. Interestingly, the centrality dependence of $\langle v_{2}\rangle$ and $\sigma_{v_{2}}$ are complementary to each other for the Woods-Saxon case in O–O collisions. However, this correlation between $\langle v_{2}\rangle$ and $\sigma_{v_{2}}$ is not completely applicable for O–O collisions with an $\alpha$-cluster nuclear density profile. Figure 7 (d) shows the evolution of relative elliptic flow fluctuation ($F(v_{2})$) as a function of collision centrality in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV. Interestingly, we find a similar centrality dependence of $F(v_{2})$ as found for $\sigma_{v_{2}}$. The centrality dependence of $F(v_{2})$ in O–O collisions with Woods-Saxon nuclear distribution resembles a p–Pb collision [93]. In contrast, for $\alpha$-cluster nuclear density profiles, $F(v_{2})$ is higher for most central collisions and decreases towards the mid-central collisions. This is an important observation where we find that a change in the initial nucleon distribution (Woods-Saxon versus $\alpha$-cluster) can cause a difference in the collision centrality dependence of relative elliptic flow fluctuations for a hydro-dynamically evolving system.
IV Summary
In this paper, we study the anisotropic flow coefficients in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using a hybrid of IPGlasma + MUSIC + iSS + UrQMD model and try to probe the final state effects of the presence of an exotic $\alpha$-cluster nuclear geometry in contrast to a uniform Woods-Saxon nuclear distribution inside a 16O nucleus. Here, we study the centrality dependence of anisotropic flow coefficients. In addition, for the first time, we report studies of elliptic flow fluctuations as a motivation to distinguish O–O collisions having $\alpha$-cluster geometry from a regular Woods-Saxon distribution. The observation of large triangular flow in the most central O–O collisions having an $\alpha$-clustered nuclei compared to a Woods-Saxon nuclear distribution is consistent with transport model predictions. In addition, we observe a significant centrality dependence of relative elliptic flow fluctuations in both the nuclear distributions. With this study, it is clear that the presence of an $\alpha$-cluster nuclear geometry has a significant effect on both elliptic and triangular flow. However, we note a distinctive feature appears when observables related to fluctuations are studied, i.e., triangular flow and elliptic flow fluctuations. While probing a nuclear density profile is a subject matter of low-energy nuclear physics, it will be very interesting to find an appropriate observable that can be measured in TeV nuclear collisions to do the job. This makes the present study more interesting.
Acknowledgements
S. P. acknowledges the doctoral fellowship from UGC, Government of India. R. S. sincerely acknowledges the DAE-DST, Government of India funding under the mega-science project – “Indian participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021-IITI (E-37123). G. G. B. gratefully acknowledges the Hungarian National Research, Development and Innovation Office (NKFIH) under Contracts No. OTKA K135515, No. NKFIH NEMZ_KI-2022-00031 and Wigner Scientific Computing Laboratory (WSCLAB, the former Wigner GPU Laboratory). The authors gratefully acknowledge the MoU between IIT Indore and Wigner Research Centre for Physics (WRCP), Hungary, for the techno-scientific international cooperation.
Appendix
.1 Centrality versus impact parameter
Figure 8: Average value of impact parameter as a function of centrality class in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD for a Woods-Saxon and $\alpha$-cluster nuclear density profile.
Figure 8 shows the average value of the impact parameter ($\langle b\rangle$) as a function of centrality class in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD for a Woods-Saxon and $\alpha$-cluster nuclear density profiles. As expected, one finds a significant dependence of $\langle b\rangle$ on collision centrality selection based on the final state-charged particle multiplicity in the V0 acceptance region, i.e., the collisions having a lower impact parameter tend to produce a higher charged particle yield at the final state. In addition, one finds a significant dependence of $\langle b\rangle$ on the nuclear density profile of the colliding oxygen nucleus to produce a similar charge particle yield in the final state. For the $\alpha$-cluster case, one finds that $\langle b\rangle$ is significantly smaller compared to a Woods-Saxon density profile because of the compact nature of $\alpha$-cluster nuclear structure, the effects of which are enhanced in the peripheral collisions. In addition, the compact nature of $\alpha$-cluster nuclear density profile is also reflected in the higher yield in most central collisions, as shown in Fig. 4.
.2 Four particle cumulants
Figure 9: Upper panel shows the second and third-order four-particle Q-cumulants ($c_{n}\{4\}$) as a function of multiplicity class (%) in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. Lower panel shows the ratio of $c_{n}\{4\}$ from Woods-Saxon to a $\alpha$-cluster nuclear density profile.
The upper panel of Fig. 9 shows the centrality dependence of four-particle cumulants ($c_{n}\{4\}$) as a function of collision centrality in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. The lower panel depicts the ratio of $c_{n}\{4\}$ from Woods-Saxon to a $\alpha$-cluster nuclear density profile. One can note that, $v_{n}\{4\}$ is related with $c_{n}\{4\}$ using Eq. (15), thus a negative value of $c_{n}\{4\}$ results in a physically acceptable value of $v_{n}\{4\}$. However, a positive value of $c_{n}\{4\}$ leads to an imaginary value of $v_{n}\{4\}$, which sometimes occur when the system is dominated with non-flow contributions [93]. In the upper panel of Fig. 9, one finds a significant difference of $c_{n}\{4\}$ between two nuclear density profiles. For an $\alpha$-cluster case, $c_{2}\{4\}$ starts with a small negative value and starts to decrease further with an increase in collision centrality, i.e., towards (10-20)centrality, attains a minimum and goes on increasing, until attains a positive value beyond (30-40)centrality class. In contrast, $c_{2}\{4\}$ for a Woods-Saxon nuclear density profile starts a minimum value for the most central case, i.e., (0-10)centrality class, and starts to rise, becomes positive beyond (30-40)centrality. Interestingly, for both the nuclear density profiles, $c_{2}\{4\}$ attains a positive value beyond a similar (30-40)centrality class, which naively implies the dominance of non-flow contribution for events with low final state multiplicity. The positive values of $c_{2}\{4\}$ on and beyond (30-40)centrality class restricts us from the calculation of four-particle elliptic flow and its fluctuations below (30-40)centrality class. On the other hand, $c_{3}\{4\}$ for both the nuclear density profiles fluctuate around 0, with a much smaller value as compared to corresponding $c_{2}\{4\}$. The positive values of $c_{3}\{4\}$ correspond to large non-flow azimuthal correlations.
.3 $p_{\rm T}$-dependence of anisotropic flow
Figure 10: Upper panel shows the elliptic flow and triangular flow using two-particle Q-cumulant method as a function of transverse momentum for different multiplicity classes in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles. The lower panel shows the ratio of $v_{n}\{2\}(p_{\rm T})$ for the Woods-Saxon density profile to $\alpha$-cluster density profile.
Figure 10 shows the transverse momentum dependence of elliptic and triangular flow calculated using the two-particle Q-cumulants method in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using IPGlasma + MUSIC + iSS + UrQMD model for Woods-Saxon and $\alpha$-cluster nuclear density profiles in the upper panel. The bottom panel shows the ratio of $v_{n}\{2\}(p_{\rm T})$ the Woods-Saxon density profile to $\alpha$-cluster density profile. One finds both $v_{2}\{2\}(p_{\rm T})$ and $v_{3}\{2\}(p_{\rm T})$ rise linearly with an increase in $p_{\rm T}$, which is a exclusive feature of hydro-based models. In addition, one observes that for the most central (0-10)class, $v_{2}\{2\}(p_{\rm T})$ and $v_{3}\{2\}(p_{\rm T})$ are closer to one another, where the separation further increases as one moves towards the mid-central collisions. This is because the contributions for $v_{2}\{2\}(p_{\rm T})$ and $v_{3}\{2\}(p_{\rm T})$ are the consequences of the interplay between the geometry of the overlap region and the eccentricity fluctuations during the collisions, etc., where the eccentricity fluctuations dominate the contributions to the anisotropic flow in the central collisions. Another interesting observation is that, in the most central collisions, $v_{n}\{2\}(p_{\rm T})$ for the $\alpha$-cluster case is larger compared to corresponding values from a Woods-Saxon distribution. This observation is consistent with the Fig 5. However, towards the mid-central collisions, $v_{n}\{2\}(p_{\rm T})$ from Woods-Saxon nuclear distribution is larger compared to an $\alpha$-cluster nuclear density profile. For the central collisions, the lower ratio plot of Woods-Saxon to an $\alpha$-cluster nuclear density profile for both elliptic and triangular flow shows a flat ratio, almost independent of $p_{\rm T}$. However, it no longer holds towards the mid-central collisions, indicating a $p_{\rm T}$ dependence of medium effects for different nuclear profiles to the anisotropic flow coefficients.
Software
In this work, we used publicly available software for data generation. The software for IPGlasma, MUSIC, iSS and UrQMD can be accessed using the web links provided in Refs. [99, 100, 101, 102].